# The conformal Laplacian and positive scalar curvature metrics on manifolds with boundary

### v4

@@ 5. Manifolds with $3$-dimensional lens space boundaries
For $p$ odd (so there is a unique spin structure), Gilkey [12, Theorem 2.5 (c)] shows that with respect to the standard round metric $g_{0}$ on $L(p,q)$
@@ 5.1. A rational coboundary for lens spaces
Here is a brief outline of the argument. The AHSS has $E^{2}_{p,q}\cong H_{p}(BG;\Omega^{\,\operatorname{spin}}_{q})$ and converges to $\Omega^{\,\operatorname{spin}}_{*}(BG)$. It is standard that for a finite group, the rational homology of $BG$ vanishes in degree greater than zero. Moreover, the Anderson-Brown-Peterson calculation of the spin cobordism group [1] (see [23, page 336] for the rational calculation) implies that $\Omega^{\,\operatorname{spin}}_{q}\otimes\mathbb{Q}=0$ when $q$ is odd. It follows that the AHSS for $\Omega^{\,\operatorname{spin}}_{*}\otimes\mathbb{Q}$ collapses immediately at the $E^{2}$ term, and so $\Omega^{\,\operatorname{spin}}_{2k+1}(BG)\otimes\mathbb{Q}=0$.
@@
The construction begins with the surface $\Sigma_{p}^{0}$ drawn (for $p=3$) in Figure 1, inspired by Figure 1 in [13]. It may described as a stack of $p$ discs, joined by $p$ half-twisted bands as shown in the figure. The surface $\Sigma_{p}^{0}$ has a $p$-fold symmetry $\tau$, given by counterclockwise rotation by angle $2\pi/p$ with fixed point set $p$ points, one on each of the discs. Furthermore $\partial\Sigma_{p}^{0}$ is $p$ circles, cyclically permuted by $\tau$. Hence the symmetry extends to an action of $\mathbb{Z}_{p}$ on the closed surface
@@
If $M(p,q)$ has a metric $g$ with $\eta_{1,g}>0$ and totally geodesic boundary, then by Cor. 3.3 we can assume that $R_{i^{*}\tilde{g}}>0$ for some $\tilde{g}$ conformal to $g$. By Lem. 5.1, $\xi_{\alpha}(L(p,q),\tilde{g})\neq 0.$ $R_{i^{*}\tilde{g}}>0$ implies $\zeta_{1,\tilde{g}}>0$, so by Thm. 4.1, we have $\eta_{1,\tilde{g}}\leq 0$ and hence $\eta_{1,g}\leq 0$, a contradiction. In particular, $g$ cannot be psc with totally geodesic boundary.
By [5, Thm. 27], a psc metric $g_{1}$ with mean convex boundary can be deformed through psc metrics to a psc metric $g$ with totally geodesic boundary. Since $\eta_{1.g}>0$, $g$ and hence $g_{1}$ cannot exist. ∎
@@ 5.2. Higher dimensions
Let $(Y,g,\alpha)$ be a $(4k-1)$-dimensional Riemannian spin manifold with $k>1$, where $\alpha$ is a representation of $\pi_{1}(Y)$ that factors through a finite group. We know $N(Y,\alpha)=\partial(M,\hat{\alpha})$ for some $N$ with $M$ spin. If $\xi_{\alpha}(Y,g)\neq 0$, then $M$ admits no metric $h$ with $\eta_{1,h}>0$ and $i^{*}h=g$ on each component of $\partial M.$ Thus the $\xi$-invariant is an obstruction to extending $g$ from $NY$ to a metric $h$ on a bounding manifold $M$ such that $h$ is psc with totally geodesic boundary.
@@
    $\xi_{\alpha}(NY,g)=\sum_{i=1}^{N}\xi_{\alpha}(Y_{i},g)$ may vanish, since the $\xi_{\alpha}(Y_{i},g)$ differ by integers. However, this cannot occur if $\xi_{\alpha}(Y_{i},g)\not\in\mathbb{Q}.$ While irrational $\xi$-invariants occur for $Y=S^{1}$, we again know of no higher dimensional examples. In particular, the examples in [6, Lemma 2.3] are all rational.

### v5

@@ 5. Manifolds with $3$-dimensional lens space boundaries
For $p$ odd (so there is a unique spin structure), Gilkey [16, Theorem 2.5 (c)] shows that with respect to the standard round metric $g_{0}$ on $L(p,q)$
@@ 5.1. A rational coboundary for lens spaces
Here is a brief outline of the argument. The AHSS has $E^{2}_{p,q}\cong H_{p}(BG;\Omega^{\,\operatorname{spin}}_{q})$ and converges to $\Omega^{\,\operatorname{spin}}_{*}(BG)$. It is standard that for a finite group, the rational homology of $BG$ vanishes in degree greater than zero. Moreover, the Anderson-Brown-Peterson calculation of the spin cobordism group [3] (see [28, page 336] for the rational calculation) implies that $\Omega^{\,\operatorname{spin}}_{q}\otimes\mathbb{Q}=0$ when $q$ is odd. It follows that the AHSS for $\Omega^{\,\operatorname{spin}}_{*}\otimes\mathbb{Q}$ collapses immediately at the $E^{2}$ term, and so $\Omega^{\,\operatorname{spin}}_{2k+1}(BG)\otimes\mathbb{Q}=0$.
@@
The construction begins with the surface $\Sigma_{p}^{0}$ drawn (for $p=3$) in Figure 1, inspired by Figure 1 in [18] and [17, §5]. It may described as a stack of $p$ discs, joined by $p$ half-twisted bands as shown in the figure. The surface $\Sigma_{p}^{0}$ has a $p$-fold symmetry $\tau$, given by counterclockwise rotation by angle $2\pi/p$ with fixed point set $p$ points, one on each of the discs. Furthermore $\partial\Sigma_{p}^{0}$ is $p$ circles, cyclically permuted by $\tau$. Hence the symmetry extends to an action of $\mathbb{Z}_{p}$ on the closed surface
@@
If $M(p,q)$ has a metric $g$ with $\eta_{1,g}>0$ and totally umbilic boundary, then by Cor. 3.3 we can assume that $R_{i^{*}\tilde{g}}>0$ for some $\tilde{g}$ conformal to $g$. By Lem. 5.1, $\xi_{\alpha}(L(p,q),\tilde{g})\neq 0.$ $R_{i^{*}\tilde{g}}>0$ implies $\zeta_{1,\tilde{g}}>0$, so by Thm. 4.1, we have $\eta_{1,\tilde{g}}\leq 0$ and hence $\eta_{1,g}\leq 0$, a contradiction. In particular, $g$ cannot be psc with totally geodesic boundary.
By [7, Thm. 27], a psc metric $g_{1}$ with mean convex boundary can be deformed through psc metrics to a psc metric $g$ with totally geodesic boundary. Since $\eta_{1.g}>0$, $g$ and hence $g_{1}$ cannot exist. ∎
Remark 5.1.
If we replace mean convex boundary with minimal boundary in the theorem, the proof follows from earlier work [1], [2, Thms. 2.1, 4.6], where it is proved that such a metric can be deformed to a product metric near the boundary.
@@ 5.2. Higher dimensions
Let $(Y,g,\alpha)$ be a $(4k-1)$-dimensional Riemannian spin manifold with $k>1$, where $\alpha$ is a representation of $\pi_{1}(Y)$ that factors through a finite group. We know $N(Y,\alpha)=\partial(M,\hat{\alpha})$ for some $N$ with $M$ spin and $\hat{\alpha}$ an extension of $\alpha$. If $\xi_{\alpha}(Y,g)\neq 0$, then $M$ admits no metric $h$ with $\eta_{1,h}>0$ and $i^{*}h=g$ on each component of $\partial M.$ Thus the $\xi$-invariant is an obstruction to extending $g$ from $NY$ to a metric $h$ on a bounding manifold $M$ such that $h$ is psc with totally geodesic boundary.
@@
    $\xi_{\alpha}(NY,g)=\sum_{i=1}^{N}\xi_{\alpha}(Y_{i},g)$ may vanish, since the $\xi_{\alpha}(Y_{i},g)$ differ by integers. However, this cannot occur if $\xi_{\alpha}(Y_{i},g)\not\in\mathbb{Q}.$ While irrational $\xi$-invariants occur for $Y=S^{1}$, we again know of no higher dimensional examples. In particular, the examples in [8, Lemma 2.3] are all rational.